OpenStudy (anonymous):

Please show all work 1. You and your friend are shopping for painting supplies. Your friend wants five bottles of paint and one set of paint brushes. You want seven bottles of paint and two sets of paint brushes. Each bottle of paint costs $3.69. If you and your friend’s total cost is $51.69, what is the cost of one set of paint brushes?

OpenStudy (anonymous):

price of all paint: 3.69x12=$44.28 51.69-44.28=$7.41 7.41/how many sets of brushes=7.41/3 one set of paint brushes=$2.47

OpenStudy (anonymous):

thank you for the answer

OpenStudy (anonymous):

yup :)